# SplineSketch: Even More Accurate Quantiles with Error Guarantees

### v1

@@ Abstract.
Space-efficient estimation of quantiles in massive datasets is a fundamental problem with numerous applications in data monitoring and analysis. While theoretical research led to optimal algorithms, such as the Greenwald-Khanna algorithm or the KLL sketch, practitioners often use other sketches that perform significantly better in practice but lack theoretical guarantees. Most notably, the widely used $t$-digest has unbounded worst-case error.
In this paper, we seek to get the best of both worlds. We present a new quantile summary, SplineSketch, for numeric data, offering near-optimal theoretical guarantees and outperforming $t$-digest by a factor of 2-20 on a range of synthetic and real-world datasets with non-skewed frequency distributions. To achieve such performance, we develop a novel approach that maintains a dynamic subdivision of the input range into buckets while fitting the input distribution using monotone cubic spline interpolation. The core challenge is implementing this method in a space-efficient manner while ensuring strong worst-case guarantees.
journalyear: 2025conference: ; ; copyright: none
@@ 1. Introduction
Approximating order statistics is one of the central sketching problems. The goal is to process a massive dataset that is possibly distributed and summarize it in a small space into a sketch that is updatable with new observations, mergeable, and provides highly accurate estimates for the median, percentiles, and their generalization, quantiles. Equivalently, we can estimate data distribution by approximating ranks, i.e., the number of input items no larger than a given element $y$. Quantile sketches have been applied, e.g., to monitoring latencies (Tene, 2015) or for detecting anomalies (Siffer et al., 2017).
@@
Streaming quantile estimation has thus received a significant attention in research community. From the theory point of view, it has been solved to a large extent. Namely, optimal algorithms are known for the uniform error, such as the Greenwald-Khanna sketch (Greenwald and Khanna, 2001; Cormode and Veselý, 2020) and KLL (Karnin et al., 2016). However, most of these theory-based quantile summaries are rarely used in practice. Instead, practitioners often use algorithms that perform significantly better on real-world datasets but have no guarantees. Most notably, $t$-digest (Dunning and Ertl, 2019) is widely used and, according to (Dunning, 2021), it has been adopted by major tech companies (reportedly Microsoft, Facebook, Google) and open-source projects (Elasticsearch, TimescaleDB, Apache Dubbo, etc.). At the same time, no matter how large the $t$-digest is, its error can be arbitrarily bad (Cormode et al., 2021).
Here, we introduce a new quantile sketch, called SplineSketch, that not only gets the best of both worlds – theoretical guarantees and very high accuracy in practice – but also significantly outperforms the state of the art, including $t$-digest (Dunning and Ertl, 2019). It processes any numerical input and requires a single parameter $k$ that determines its space usage, which is $O(k)$ memory words for processing the input and $16\cdot k$ bytes when serialized for storage. The sketch deterministically guarantees a theoretically near-optimal error, close to the best possible error of $O(n/k)$ with $k$ memory words; namely, the error is $O(n/k\cdot\log\alpha)$, where $\alpha$ is the aspect ratio, defined as the ratio of the largest and the smallest difference between distinct items in the dataset (e.g., for fixed-point numbers, $\alpha$ is bounded by the universe size). At the same time, the accuracy of our sketch in practice is in fact typically 10–200 better than $n/k$. Furthermore, we prove our sketch is fully mergeable while retaining the guarantees, i.e., we analyze it in the most general setting when the sketch is constructed by an arbitrary sequence of pairwise merge operations.
@@
Figure 1. Average error of SplineSketch, KLL, MomentSketch, and $t$-digest on three real-world datasets (Sec. 5).
@@ 1.1. Overview of our approach
We develop a new technique for designing quantile sketches that draws inspiration from both theoretical and practical approaches. The core of our approach is in maintaining a subdivision of the input range into buckets in a way that better adapts to the input distribution. Namely, SplineSketch consists of thresholds $\tau_{1},\cdots,\tau_{k}$ and then $(\tau_{i-1},\tau_{i}]$ is the $i$-th bucket, with $\tau_{0}=-\infty$. For each bucket, we store a counter $b_{i}$ that represents an approximate number of items that lie in $(\tau_{i-1},\tau_{i}]$. Furthermore, we use cubic spline interpolation in order to answer queries more accurately in practice.
Figure 2. SplineSketch illustration, with the values and thresholds $\tau_{i}$ on the $x$ axis and the rank space $[0,n]$ on the $y$ axis. The true cumulative distribution function of the data is depicted as a red dashed curve. The green dots correspond to the prefix sums of bucket counters, i.e., the counter $b_{i}$ equals the difference of the $y$ values of the the green dots at $\tau_{i}$ and $\tau_{i-1}$. The solid black curve is an interpolation over the green dots.
@@
The estimated rank of any threshold $\tau_{i}$ is $\sum_{j=1}^{i}b_{i}$. It is however not clear how to answer rank queries that are not at one of the bucket boundaries $\tau_{i}$. Most of the previous work would just instead query the closest value $\tau_{j}$ and return that as the answer, or perform a linear interpolation. However, we seek to do better than that. To this end, we use cubic interpolation. Specifically, if we are querying $x$, such that $\tau_{i}<x<\tau_{i+1}$, we estimate the ranks $\hat{r}_{i-1},\cdots,\hat{r}_{i+2}$ of $\tau_{i-1},\cdots,\tau_{i+2}$. We then use a piecewise cubic Hermite interpolating polynomial (PCHIP) in order to interpolate between $\hat{r}_{i}$ and $\hat{r}_{i+1}$. Namely, we evaluate this interpolating polynomial at $x$ and return it as our estimate.
@@
The most important choice we need to make is what the buckets should be. In general, we will set $\tau_{0}$ and $\tau_{k}$ to be the smallest and the largest item, respectively, that we have seen so far. The main part is adjusting the buckets by splitting some buckets and joining pairs of adjacent buckets. When we join two adjacent buckets, we set the new counter to the sum of the two buckets’ counters. In the split procedure, we split the bucket in the middle. We then need to choose how to split the count into the two new sub-buckets. To this end, we perform a query at the midpoint (as described above) and we divide the count so as to make the value at the midpoint as returned by the query. It remains to specify how we choose which buckets to join and split. When describing this here, we ignore the question of efficiency. In the end, we will have a buffer and perform updates in batches, allowing for low amortized update time.
@@
As an invariant, to ensure we always have $k$ buckets, whenever we split a bucket, we join two buckets, which we choose as described below. There are two rules to choose a bucket to split: First, throughout the execution, we ensure that all buckets have counter $O(n/k)$, and we split a bucket when it gets too big. We use these bucket size bounds, together with other invariants, to show the worst-case error bounds. Intuitively speaking, we show that no item can be part of many splits and that each split can only contribute to the error by at most the size of the bucket the item belongs to.
@@
Second, while we want to get good error bounds in the worst case, we seek to perform significantly better in practice. To this end, we also introduce for each bucket a number which we call heuristic error, which is our guess for the actual error of the cubic spline inside the bucket. The underlying reason is that how the data distribution function changes is captured by the derivatives of the distribution. For example, suppose that the data items are i.i.d. from an unknown distribution. If the pdf $f$ does not change too much around a bucket, i.e., the pdf’s derivative $f^{\prime}$ is small, the error should be low as well. Of course, we do not know $f$ and data may not be i.i.d., but an estimate of the pdf’s derivative around a bucket serves as a heuristic to the true error. This number is not stored but we compute it when we need it. If the maximum heuristic error can be decreased substantially by splitting one bucket and joining two other adjacent buckets, then we perform the split and join of these buckets. The exact function that we use as the heuristic error is discussed in Section 3.1 (see Equantion 2).
When we need to perform a join (typically when we split a bucket), we would like to join two adjacent buckets whose joining would result in a bucket with the lowest possible value of heuristic error. However, this may not be possible for two reasons. First, it may violate the bucket size bound – this can be easily fixed by only considering those pairs of buckets where the bucket size bound will not be violated after the join. In fact, we do not join two buckets if the resulting bucket would be, say, 75% full. Second, in some situations this may lead to too many joins and splits on a single bucket, which would invalidate our worst-case bounds. To this end, we divide the input into epochs, where one epoch is defined as time during which the input size increased by a fixed constant factor. We then keep a bitmap to ensure that a bucket that was created by a split will not be joined during the same epoch, that is, we protect thresholds of newly created buckets from joining for a certain time period. This allows us to prove the worst-case bounds.
@@
The merge operation of our sketch is conceptually straightforward: First, we estimate the ranks of thresholds of each sketch in the other sketch. Then we take the union of the two sets of thresholds, disregarding duplicate thresholds or thresholds that are too close to another threshold, which results in a sketch with possibly too many buckets. We set the counters of new buckets based on their estimated ranks in both sketches. Finally, we join enough pairs of adjacent buckets based on their heuristic error so that the resulting sketch has the desired number of buckets. To ensure worst-case guarantees, the protection of thresholds from joining is inherited from the larger sketch if merging two sketches of substantially different size.
@@
Using joining or splitting the buckets allows us to easily change the number $k$ of buckets. Namely, when we want to make the sketch smaller, we just perform the appropriate number of buckets joins without any splits, whereas for increasing the sketch size, we perform splits without joins. This demonstrates the flexibility of our approach, and is useful in practice, e.g., when we allocate more memory for processing the input and then compress the sketch into much smaller size.
1.2. Related work
@@
Quantile summaries are intensively investigated since the 90s, with the pioneering work of Munro and Paterson from 1978 (Munro and Paterson, 1980) who showed that exact quantile selection in sublinear space is impossible. In the comparison-based model, Greenwald and Khanna (Greenwald and Khanna, 2001) designed a deterministic algorithm that stores $O(\varepsilon^{-1}\cdot\log\varepsilon n)$ stream items and guarantees ±$\pm\varepsilon\cdot n$ uniform error; this bound is optimal among deterministic comparison-based quantile summaries (Cormode and Veselý, 2020). However, the GK summary is not known to be mergeable while retaining its space bounds (Agarwal et al., 2013), despite that the intricate analysis from (Greenwald and Khanna, 2001) was recently simplified and generalized to weighted updates (Assadi et al., 2023; Gribelyuk et al., 2024). Randomization allows to remove the logarithmic dependency on the stream length $n$, and after a sequence of improvements (Manku et al., 1999; Luo et al., 2016; Felber and Ostrovsky, 2015), Karnin, Lang, and Liberty (Karnin et al., 2016) developed an optimal randomized algorithm that stores just $O(\varepsilon^{-1})$ items (with a constant probability of a too large error). Beyond the comparison-based model, $q$-digest (Shrivastava et al., 2004) builds an implicit binary tree over the items’ universe $\mathcal{U}$ and stores a subset of the tree nodes. It requires space of $O(\varepsilon^{-1}\cdot\log|\mathcal{U}|)$ and a foreknowledge of $\mathcal{U}$. Very recently, the space usage of $q$-digest was decreased to the optimal $O(1/\varepsilon)$ memory words, by packing more nodes and their associated counters into one memory word (Gupta et al., 2024).
@@
Another line of work focused on the stronger relative (multiplicative) error guarantee, which requires that the error for an item of rank $r$ is at most ±$\pm\varepsilon\cdot r$. For deterministic comparison-based algorithms, there is a gap of $O(\log\varepsilon n)$ between the merge-and-prune algorithm using space $O(\varepsilon^{-1}\cdot\log^{3}\varepsilon n)$ (Zhang and Wang, 2007) and the lower bound of $\Omega(\varepsilon^{-1}\cdot\log^{2}\varepsilon n)$ (Cormode and Veselý, 2020). Efficient randomized sketches with relative error have only appeared recently, namely, ReqSketch with space $O(\varepsilon^{-1}\cdot\log^{1.5}\varepsilon n)$ (Cormode et al., 2023) and its “elastic“ version achieving space close the information-theoretic lower bound of $\Omega(\varepsilon^{-1}\cdot\log\varepsilon n)$ up to asymptotically smaller factors (Gribelyuk et al., 2025); however, mergeability of the latter sketch is open. There is also an extension of $q$-digest to the relative error that uses space $O(\varepsilon^{-1}\cdot\log\varepsilon n\cdot\log|\mathcal{U}|)$ (Cormode et al., 2006); it is not known to be optimal.
@@
For dynamic streams (i.e., with deletions), one can only achieve the uniform error guarantee in sublinear space, and the best quantile summary supporting deletions is the Dyadic CountSketch (Luo et al., 2016).
@@
Finally, there are more heuristic approaches that do not aim for any worst-case guarantees and work only for numerical data, similarly as our algorithm. Specifically, $t$-digest (Dunning and Ertl, 2019; Dunning, 2021) performs an online one-dimensional $k$-means clustering by averaging data items into a given number of centroids, and interpolates linearly between the centroids. A downside of this approach is that many of the items summarized by a centroid may be smaller or larger than an adjacent centroid mean, leading under- or over-estimation. That is, the centroids are only ordered by their means, while the represented items may be far from ordered. In fact, one can impose high overlaps of represented items, leading to almost arbitrarily large error on adversarial instances (Cormode et al., 2021). Linear interpolation techniques were combined with the KLL sketch to improve its accuracy (Schiefer et al., 2023). In contrast to rank error, DDSketch (Masson et al., 2019) and UDDSketch (Epicoco et al., 2020) provide the relative value error (i.e., an item close to the desired quantile in the item space) based on maintaining a suitable exponential histogram. Finally, MomentSketch (Gan et al., 2018; Mitchell et al., 2021) is possibly the most compact summary as it consists of $k$ moments and log-moments of data items, for a small $k$ such as $k=15$. Upon a query, the sketch constructs a distribution with the same moments and log-moments, using the maximum entropy principle and Chebyshev polynomials.
2. Preliminaries
Throughout the paper, $n$ is the number of items currently summarized by the sketch. For a fixed multiset $S$ of real numbers and for any $x\in\mathbb{R}$ let $\mathrm{rank}(x)=|\{y\in S:y\leq x\}|$, i.e., the number of input items smaller or equal to $x$. For any $q\in(0,1]$, we call $y:=\min\{x:\mathrm{rank}(x)\geq q\cdot n\}$ the $q$-quantile of $S$. This definition implies that a method for estimating ranks may be used for estimating quantiles with the same accuracy, by performing binary search. The aspect ratio $\alpha$ of an input $x_{1},\dots,x_{n}$ is defined as $\alpha=(\max_{i}x_{i}-\min_{i}x_{i})/\gamma$, where $\gamma=\min_{i\neq j:x_{i}\neq x_{j}}|x_{i}-x_{j}|$ is the smallest difference of distinct items on input (assuming at least two distinct items are present).
We use a monotone variant of the Piecewise Cubic Hermite Interpolating Polynomial (PCHIP) (Fritsch and Butland, 1984; Fritsch and Carlson, 1980), which is an interpolation method that constructs a continuous function by fitting cubic polynomials between data points. Unlike standard cubic spline interpolation, PCHIP adjusts the first derivatives at each data point to preserve the shape and monotonicity of the original data. For an introduction at a greater detail, including the formulas for the interpolation, see (Wong, 2020). Since the ranks of the input items are non-decreasing, preserving monotonicity is an important feature of the interpolation, rendering PCHIP suitable for our application.

### v2

@@ Abstract.
Space-efficient streaming estimation of quantiles in massive datasets is a fundamental problem with numerous applications in data monitoring and analysis. While theoretical research led to optimal algorithms, such as the Greenwald-Khanna algorithm or the KLL sketch, practitioners often use other sketches that perform significantly better in practice but lack theoretical guarantees. Most notably, the widely used $t$-digest has unbounded worst-case error.
In this paper, we seek to get the best of both worlds. We present a new quantile summary, SplineSketch, for numeric data, offering near-optimal theoretical guarantees, namely uniformly bounded rank error, and outperforming $t$-digest by a factor of 2-20 on a range of synthetic and real-world datasets. To achieve such performance, we develop a novel approach that maintains a dynamic subdivision of the input range into buckets while fitting the input distribution using monotone cubic spline interpolation. The core challenge is implementing this method in a space-efficient manner while ensuring strong worst-case guarantees.
quantile estimation, streaming algorithms, mergeable summaries, data sketches, cubic spline interpolation
journalyear: 2025conference: ; ; copyright: noneccs: Theory of computation Sketching and samplingccs: Theory of computation Streaming modelsccs: Theory of computation Data structures and algorithms for data management
@@ 1. Introduction
Approximating order statistics is a central sketching problem. The goal is to process a massive dataset that is possibly distributed and summarize it in a small space into a sketch that is updatable with new observations, mergeable, and provides highly accurate estimates for the median, percentiles, and their generalization, quantiles. Equivalently, we can estimate the data distribution by approximating ranks, i.e., the number of input items no larger than a given element. Quantile sketches have been applied, e.g., to monitoring latencies (Tene, 2015) or for detecting anomalies (Siffer et al., 2017).
@@
Streaming quantile estimation has thus received a significant attention in research community. From the theory point of view, it has been solved to a large extent. Namely, optimal algorithms are known for the uniform error, such as the Greenwald-Khanna (GK) sketch (Greenwald and Khanna, 2001; Cormode and Veselý, 2020) and KLL (Karnin et al., 2016). However, the accuracy of these theory-based quantile summaries on real-world data is only a small constant factor better than the worst-case bound. Instead, practitioners often use algorithms that perform significantly better in practice but have no guarantees. Most notably, $t$-digest (Dunning and Ertl, 2019) is widely used and, according to (Dunning, 2021), it has been adopted by major tech companies (reportedly Microsoft, Facebook, Google) and open-source projects (Elasticsearch, TimescaleDB, Apache Dubbo, etc.). At the same time, no matter how large the $t$-digest is, its error can be arbitrarily bad (Cormode et al., 2021). Furthermore, as we outline in Sec. 2, to the best of our knowledge, other quantile sketches do not provide both worst-case bounds and error similar to $t$-digest in practice.
Here, we introduce a new quantile sketch, called SplineSketch, that not only gets the best of both worlds – theoretical guarantees and very high accuracy in practice – but also significantly outperforms the state of the art, including $t$-digest (Dunning and Ertl, 2019). Our sketch processes any numerical input in a streaming fashion and is fully mergeable, enabling efficient parallel or distributed processing of large datasets. Before explaining our contribution in more detail, we define our setting, including the considered error guarantees.
Definitions.
We use $n$ for the number of items currently summarized by the sketch. For a multiset $S$ of real numbers and any $x\in\mathbb{R}$, let $\mathrm{rank}(x)=|\{y\in S:y\leq x\}|$, i.e., the number of input items smaller than or equal to $x$. For any $q\in(0,1]$, we call $y:=\min\{x\in S:\mathrm{rank}(x)\geq q\cdot n\}$ the $q$-quantile of $S$, i.e., the $q$-quantile is the $\lceil q\cdot n\rceil$-th largest item in $S$ (if there are no duplicates). This definition implies that a method for estimating ranks may be used for estimating quantiles with the same accuracy, by performing binary search. The aspect ratio $\alpha$ of $S$ is defined as $\alpha=(\max(S)-\min(S))/\gamma$, where $\gamma=\min_{i\neq j:x_{i}\neq x_{j}}|x_{i}-x_{j}|$ is the smallest difference of distinct items on input, assuming at least two distinct items are present.
We focus on the streaming and mergeability settings. The streaming model requires that the dataset $S$ is processed in one pass with low memory (in an arbitrary order, possibly adversarial). In the more general setting of mergeable summaries (Agarwal et al., 2013), $S$ is arbitrarily split into a number of parts, each summarized in the streaming fashion and the resulting summaries are combined via pairwise merge operations in an arbitrary way. In other words, besides answering rank and quantile queries, the sketch needs to have an update operation, which adds a new observation to the dataset, and a merge operation, which takes two sketches and outputs a single sketch of their combined input.
Following a long line of work, e.g., (Shrivastava et al., 2004; Greenwald and Khanna, 2001; Agarwal et al., 2013; Karnin et al., 2016; Luo et al., 2016; Cormode and Veselý, 2020; Assadi et al., 2023; Gupta et al., 2024; Gribelyuk et al., 2024), we study the uniform rank error which for a given accuracy parameter $\varepsilon>0$ and any query item $x$ (that may or may not be in the input) requires the algorithm to return a rank estimate $\widehat{\mathrm{rank}}(y)$ such that $|\widehat{\mathrm{rank}}(x)-\mathrm{rank}(x)|\leq\varepsilon\cdot n$, and for any quantile query $q\in(0,1]$, the algorithm should return a $q^{\prime}$-quantile for $q^{\prime}=q\pm\varepsilon$.
@@
Figure 1. Average error of SplineSketch, KLL, MomentSketch, $t$-digest, and GKAdaptive on three real-world datasets (Sec. 5).
Our contributions
Our main contribution is a new technique for designing quantile summaries with bounded worst-case error and high accuracy in practice, combining dynamically maintained subdivision of the input range with a suitable interpolation over the buckets, outlined in Section 1.1 and described in Section 3.
We demonstrate the power of this technique on a new quantile sketch, called SplineSketch. It requires a single parameter $k$ that determines its space usage, which is $O(k)$ memory words for processing the input and $16\cdot k$ bytes when serialized for storage. The sketch deterministically guarantees a theoretically near-optimal uniform rank error on any input, close to the best possible error of $O(n/k)$ with $k$ memory words; namely, the error is $O(n/k\cdot\log\alpha)$, where $\alpha$ is the aspect ratio. At the same time, the accuracy of our sketch in practice is in fact typically 10–200 times better than $n/k$. Furthermore, we prove our sketch is fully mergeable while retaining the guarantees, i.e., we analyze it in the most general setting when the sketch is constructed by an arbitrary sequence of pairwise merge operations. Finally, our sketch can be resized based on the current memory requirements.
@@ 1.1. Overview of our approach
The core of our approach is in maintaining a subdivision of the input range into buckets in a way that better adapts to the input distribution. Namely, SplineSketch consists of thresholds $\tau_{1},\cdots,\tau_{k}$ and then $(\tau_{i-1},\tau_{i}]$ is the $i$-th bucket, with $\tau_{0}=$-$\infty$. For each bucket, we store a counter $b_{i}$ that represents an approximate number of items that lie in $(\tau_{i-1},\tau_{i}]$; we refer to $b_{i}$ as the size of the bucket $i$. Furthermore, we use cubic spline interpolation in order to answer queries more accurately in practice; see Fig. 2 for an illustration.
Figure 2. SplineSketch illustration, with the values and thresholds $\tau_{i}$ on the $x$ axis and the rank space $[0,n]$ on the $y$ axis. The true cumulative distribution function of the data is depicted as a red dashed curve. The green dots correspond to the prefix sums of bucket counters, i.e., the counter $b_{i}$ is the difference of the $y$ values of the green dots at $\tau_{i}$ and $\tau_{i-1}$. The solid black curve is an interpolation over the green dots.
@@
The estimated rank of any threshold $\tau_{i}$ is $\sum_{j=1}^{i}b_{i}$. It is however not clear how to answer rank queries that are not at one of the bucket boundaries $\tau_{i}$. Most of the previous work would just instead query the closest value $\tau_{j}$ and return that as the answer, or perform a linear interpolation. However, we seek to do better than that, using more advanced interpolation methods, namely the piecewise cubic Hermite interpolating polynomial (PCHIP) interpolation (Fritsch and Butland, 1984; Fritsch and Carlson, 1980). Specifically, if we are querying $x$, such that $\tau_{i}<x<\tau_{i+1}$, we estimate the ranks $\hat{r}_{i-1},\cdots,\hat{r}_{i+2}$ of $\tau_{i-1},\cdots,\tau_{i+2}$, then compute the PCHIP interpolation in order to interpolate between $\hat{r}_{i}$ and $\hat{r}_{i+1}$, and evaluate this polynomial at $x$, which is our estimate.
@@
The most important choice we need to make is what the thresholds of the buckets should be. We always set $\tau_{1}$ and $\tau_{k}$ to be the smallest and the largest item, respectively, that we have seen so far. Other than that, as the new data keeps arriving, we will be adjusting the buckets by splitting some of them and joining some pairs of adjacent ones. When we join two adjacent buckets, we simply remove the threshold that they share and set the new counter to the sum of the two buckets’ counters. In the split procedure, we split the bucket in the middle. We then need to choose how to split the count into the two new sub-buckets. To this end, we perform a query at the midpoint (as described above) and we divide the count so as to make the value at the midpoint as returned by the query. It remains to specify how we choose which buckets to split and join. In the description below, we ignore the efficiency. In the end, we use a buffer and perform updates in batches, obtaining low amortized update time.
@@
There are two rules for choosing a bucket to split. Firstly, throughout the execution, we ensure that all buckets have counter $O(n/k)$, and we split a bucket when it gets too big (the exact value of the bucket bound is defined in Sec. 3.1). We later use these size bounds, together with other invariants, to show the worst-case error bounds. Intuitively speaking, we show that no item can be part of too many splits and that each split can only contribute to the error by at most the size of the bucket the item belongs to.
@@
The second rule for bucket splitting is intended to get significantly better error in practice, while preserving the good error bounds in the worst case. To this end, we also introduce for each bucket a number which we call heuristic error, which is our guess for the actual error of the cubic spline inside the bucket. The underlying reason is that how the data distribution function changes is captured by the derivatives of the distribution. To illustrate the intuition behind this notion, suppose the data items are i.i.d. from an unknown distribution. If the growth of its CDF $F$ does not change too much around a bucket, specifically the absolute value of its second derivative $F^{\prime\prime}$ is small, the error of the spline approximation should be low as well. In practice, we do not know the CDF and the data may not be i.i.d., but we nevertheless use an estimate of the CDF’s second derivative around a bucket as a heuristic approximation of the true error. We do not store this number and instead compute it when needed. If the maximum heuristic error can be decreased substantially by splitting one bucket and joining two other adjacent buckets, while preserving bucket bounds, then we perform the split and join of these buckets. The exact function that we use as the heuristic error is in Eqn. 2 in Sec. 3.1.
To ensure we always have exactly $k$ buckets, whenever we split a bucket, we join some pair of adjacent buckets. When we perform a join, we would like to choose two adjacent buckets which together would form a bucket with the lowest possible value of the heuristic error. However, we need to ensure that the new bucket would not grow too large. Therefore, we only consider pairs for which the size of the resulting bucket would be not too close to the bucket bound, specifically at most 75% of the bucket bound.
Finally, we do not want to perform too many joins and splits on a single bucket, as it causes the error to grow. To this end, we divide the input into epochs, where one epoch is defined as time during which the input size increased by a fixed constant factor. If the bucket was created by a split, we mark its thresholds as protected and we do not allow it to be joined during the same epoch. We prove that despite all these constraints after a split there will be always at least one pair of adjacent buckets that can be joined.
@@
The merge operation for our sketch is conceptually simple: Given two SplineSketches on input, we first create a combined set of thresholds by taking the union of both sketches’ thresholds and removing duplicates. We then assign counters to the new buckets based on the sum of their estimated ranks in both original sketches, for which we use precomputed interpolations over the input sketches. Taking the union of thresholds typically results in a sketch with more buckets than desired, so we reduce the number of buckets by joining adjacent ones with the lowest heuristic error, until we reach the target bucket count. To maintain worst-case guarantees, if the two sketches being merged differ significantly in size, the threshold protection from the larger sketch is preserved during merging; however, when their size is similar, we basically reset the epoch, together with the threshold protection.
@@
Using joining or splitting the buckets allows us to easily change the number $k$ of buckets. Namely, when we want to make the sketch smaller, we just join an appropriate number of adjacent buckets without any splits, whereas for increasing the sketch size, we perform splits without joins. This demonstrates the flexibility of our approach, and can be useful in practice, e.g., when we allocate more memory for processing the input and then compress the sketch into much smaller size for storage, which leads to overall better accuracy than using the smaller size for processing.
Paper outline.
Section 2 outlines the most relevant work on quantile sketches and provides a comparison with our approach. The full algorithmic description of SplineSketch appears in Section 3, together with its theoretical properties. The implementation details appear in Section 4 and the experimental evaluation in Section 5. We conclude with discussing the pros and cons of our approach and the open problems in Section 6.
2. Related work
@@
Quantile summaries are intensively investigated since the 90s, with the pioneering work of Munro and Paterson from 1978 (Munro and Paterson, 1980) who showed that exact quantile selection in the streaming setting, i.e., one pass with sublinear space, is impossible. In the comparison-based model, Greenwald and Khanna (Greenwald and Khanna, 2001) designed a deterministic algorithm that stores $O(\varepsilon^{-1}\cdot\log\varepsilon n)$ stream items and guarantees $\pm\varepsilon\cdot n$ uniform error; this bound is optimal among deterministic comparison-based quantile summaries (Cormode and Veselý, 2020). However, the GK summary is not known to be mergeable while retaining its space bounds (Agarwal et al., 2013), despite that the intricate analysis from (Greenwald and Khanna, 2001) was recently simplified and generalized to weighted updates (Assadi et al., 2023; Gribelyuk et al., 2024). Randomization allows to remove the logarithmic dependency on the stream length $n$, and after a sequence of improvements (Manku et al., 1999; Luo et al., 2016; Felber and Ostrovsky, 2015), Karnin, Lang, and Liberty (Karnin et al., 2016) developed an optimal randomized algorithm that stores just $O(\varepsilon^{-1})$ items (with a constant probability of a too large error). Beyond the comparison-based model, $q$-digest (Shrivastava et al., 2004) uses an implicit binary tree over the items’ universe $\mathcal{U}$ and stores a subset of the tree nodes with associated counters. It requires space of $O(\varepsilon^{-1}\cdot\log|\mathcal{U}|)$ and a foreknowledge of $\mathcal{U}$. Very recently, Gupta, Singhal, and Wu developed a compressed variant of $q$-digest that fits into $O(1/\varepsilon)$ memory words, by packing more nodes into one memory word (Gupta et al., 2024). However, unlike SplineSketch, all of these algorithms do not offer practical accuracy on real-world data that goes beyond their worst-case behavior, as demonstrated in Section 5.
@@
Another line of work focused on the stronger relative (multiplicative) error guarantee, which requires that the error for an item of rank $r$ is at most $\pm\varepsilon\cdot r$. For deterministic comparison-based algorithms, there is a gap of $O(\log\varepsilon n)$ between the merge-and-prune algorithm using space $O(\varepsilon^{-1}\cdot\log^{3}\varepsilon n)$ (Zhang and Wang, 2007) and the lower bound of $\Omega(\varepsilon^{-1}\cdot\log^{2}\varepsilon n)$ (Cormode and Veselý, 2020). Efficient randomized sketches with relative error have only appeared recently, namely, ReqSketch with space $O(\varepsilon^{-1}\cdot\log^{1.5}\varepsilon n)$ (Cormode et al., 2023) and its “elastic“ version achieving space close the information-theoretic lower bound of $\Omega(\varepsilon^{-1}\cdot\log\varepsilon n)$ (Gribelyuk et al., 2025); however, mergeability of the “elastic” version is open, while ReqSketch is analyzed also in the mergeability setting. There is also an extension of $q$-digest to the relative error that uses space $O(\varepsilon^{-1}\cdot\log\varepsilon n\cdot\log|\mathcal{U}|)$ (Cormode et al., 2006); it is not known to be optimal. Compared to SplineSketch, the main benefit of relative error is better accuracy for the tails which, however, comes at the cost of higher space complexity, both in theory (additional dependency on $\log n$) and practice (see e.g. (Cormode et al., 2021)). Thus, relative-error sketches are incomparable to SplineSketch and other uniform-error algorithms. Finally, we note that even stronger error guarantees, such as the q-error as defined in (Moerkotte et al., 2014), cannot be achieved in the streaming or mergeability settings with sketches of sublinear size, due to standard space lower bounds from the one-way communication complexity of indexing.
@@
For dynamic streams (i.e., with deletions), one can only achieve the uniform error guarantee in sublinear space, and the best quantile summary supporting deletions is Dyadic CountSketch (Luo et al., 2016). The support for deletions comes at an additional cost of a polylogarithmic dependency on the universe size $|\mathcal{U}|$, while insertion-only sketches, including KLL or SplineSketch, are much more space-efficient.
In contrast to rank error, DDSketch (Masson et al., 2019) and UDDSketch (Epicoco et al., 2020) provide the relative value error (i.e., an item close to the desired quantile in the item space) based on maintaining a suitable exponential histogram. While this usually captures the distribution tails well, the rank error can be arbitrarily large as these sketches in fact fail to capture most of the distribution in sufficient detail, leading to overall worse accuracy than KLL, $t$-digest, or SplineSketch. Furthermore, the value error is not invariant to natural data transformations such as translations.
@@
Finally, there are more heuristic approaches that, similarly as SplineSketch, work only for numerical data, but do not aim for any worst-case rank guarantees, unlike SplineSketch. Specifically, $t$-digest (Dunning and Ertl, 2019; Dunning, 2021) performs an online one-dimensional $k$-means clustering by averaging data items into a given number of centroids, and interpolates linearly between the centroids. A downside of this approach is that many of the items summarized by a centroid may be smaller or larger than an adjacent centroid mean, implying under- or over-estimation. That is, the centroids are only ordered by their means, while the represented items may be far from ordered. In fact, one can impose high overlaps of represented items, leading to almost arbitrarily large error on adversarial instances (Cormode et al., 2021). Linear interpolation techniques were combined with the KLL sketch to improve its accuracy (Schiefer et al., 2023).
Finally, MomentSketch (Gan et al., 2018; Mitchell et al., 2021) is possibly the most compact summary as it consists of $k$ moments and log-moments of data items, for a small $k$ such as $k=15$. Upon a query, the sketch constructs a distribution with the same moments and log-moments, using the maximum entropy principle and Chebyshev polynomials, which is a costly operation. However, MomentSketch only works well for certain smooth distributions but, as our experiments also demonstrate, suffers a large error in many real-world cases, including the uniform distribution.
The aforementioned quantile sketches form a basis for per-key quantile estimation (a simultaneous estimation of quantiles for multiple substreams in the main stream), as investigated in a recent line of work (Dong et al., 2024; Guo et al., 2023; He et al., 2023; Shahout et al., 2023).
Additionally, we remark that the idea of approximating the data distribution using splines is not new and has been widely used in computer science before, e.g., for query optimization (Neumann and Michel, 2008) and indexing (Kipf et al., 2020) in database systems.